### 04 Jan 2024 v2

# Does the Hamiltonian determine the tensor product structure and the 3d space?

Abstract
It was proposed that the tensor product structure of the Hilbert space is uniquely determined by the Hamiltonian’s spectrum, for most finite-dimensional cases satisfying certain conditions.
I show that any such method would lead to infinitely many tensor product structures. The dimension of the space of solutions grows exponentially with the number of qudits. In addition, even if the result were unique, such a Hamiltonian would not entangle subsystems.
These results affect some proposals to recover the 3d space from the Hamiltonian.
tensor product structure; entanglement; emergent spacetime; Hilbert space fundamentalism.
I Introduction
The usual ingredients for specifying a quantum theory are a Hilbert space $\mathcal{H}$, a time-dependent density operator $\widehat{\mathbf{\rho}}(t)$ on $\mathcal{H}$ representing the total state of the system, a Hermitian operator $\widehat{\mathbf{H}}$ on $\mathcal{H}$ – the Hamiltonian operator – encoding the time evolution of the system, a decomposition of the Hilbert space as a tensor product
$\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\ldots,$ (1)
describing the structure of the system in terms of subsystems or regions of space, and a set of Hermitian operators on $\mathcal{H}$ or on the subsystem spaces $\mathcal{H}_{j}$, to represent physically observable properties like positions and momenta.
The decomposition (1) represents the tensor product structure (TPS) of $\mathcal{H}$. If the system consists of subsystems, for example particles whose number remains constant, the Hilbert spaces $\mathcal{H}_{j}$ correspond to these subsystems. Decompositions like (1) also appear when we are interested in the quantum fields on regions of space.
In practice, physicists inferred the TPS from experiments, that is, from the observables corresponding to subsystems. The TPS can be determined uniquely by a set of algebras of observables [11].
In [3] it was proposed that under certain conditions there is a method determining, only from the Hamiltonian’s spectrum (including the multiplicities of eigenvalues), a TPS supposed to be unique in most cases with $\dim\mathcal{H}<\infty$. All this without using other structures, observables, or a preferred basis. This suggestion was used in the literature to propose that the $3$d space or spacetime emerges from the Hamiltonian [2, 1, 6, 10]. In [8] I showed that if such a method results in a unique TPS, that TPS can’t have observable consequences.
In Section §II I show that no method to obtain a TPS only from the spectrum of the Hamiltonian can have a unique result. In Section §III I show that even if the resulting TPS were unique, such a Hamiltonian would be unable to entangle subsystems. So even if “unique” would mean “unique up to an equivalence” in whatever sense, such a TPS won’t be like the TPSs known from physics. I discuss the implications for some approaches to emergent space or spacetime.
II Proof of non-uniqueness
Theorem 1.
No method to obtain the TPS from the Hamiltonian’s spectrum can give a unique result. The number of additional continuous parameters needed to fix the TPS uniquely is exponential in the number of qudits.
Proof.
Any TPS obtained uniquely only from the Hamiltonian’s spectrum should be preserved by any unitary symmetry transformation $\widehat{\mathbf{S}}$ preserving the Hamiltonian, $\widehat{\mathbf{S}}\widehat{\mathbf{H}}=\widehat{\mathbf{H}}\widehat{\mathbf{S}}$. Otherwise, the method to obtain the TPS is based on additional assumptions, introducing more information than the Hamiltonian’s spectrum. To see this, consider that the method to obtain the TPS consists of calculations done in a basis of $\mathcal{H}$ in which $\widehat{\mathbf{H}}$ is diagonal, with the eigenvalues sorted in ascending order. If the resulting TPS depends only on the spectrum of $\widehat{\mathbf{H}}$, we should obtain the same TPS if we use a different basis in which $\widehat{\mathbf{H}}$ has the same form. Any such basis can be obtained from the former basis by a unitary transformation that commutes with $\widehat{\mathbf{H}}$.
The TPS can be specified uniquely in terms of algebras of observables $\mathcal{A}_{1},\mathcal{A}_{2},\ldots$, i.e. algebras of operators on $\mathcal{H}$, required to generate the full algebra of observables of $\mathcal{H}$, and to satisfy the conditions [11]
$\left[\mathcal{A}_{j},\mathcal{A}_{k}\right]=\widehat{\mathbf{0}}\textnormal{ and }\mathcal{A}_{j}\cap\mathcal{A}_{k}=\mathbb{C}\widehat{\mathbf{1}}_{\mathcal{H}}\textnormal{ for all }j\neq k.$ (2)
Therefore, the TPS resulting by transforming the original TPS using a unitary transformation $\widehat{\mathbf{S}}$ is specified by the algebra of observables
$\big{(}\mathcal{A}^{\prime}_{j}\big{)}_{j\in\{1,\ldots,n\}}=\big{(}\widehat{\mathbf{S}}\mathcal{A}_{j}\widehat{\mathbf{S}}^{\dagger}\big{)}_{j\in\{1,\ldots,n\}}.$ (3)
If the TPS were unique, this would imply that, for any $\widehat{\mathbf{S}}$ satisfying $\widehat{\mathbf{S}}\widehat{\mathbf{H}}=\widehat{\mathbf{H}}\widehat{\mathbf{S}}$, the two sets of algebras of observables are identical,
$\left\{\mathcal{A}^{\prime}_{j};j=1,\ldots,n\right\}=\left\{\widehat{\mathbf{S}}\mathcal{A}_{j}\widehat{\mathbf{S}}^{\dagger};j=1,\ldots,n\right\}.$ (4)
Therefore, a necessary condition for the uniqueness of the TPS is that the symmetries preserving the Hamiltonian should also preserve the TPS. That is,
$\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}\subseteq\textnormal{U}_{\textnormal{TPS}},$ (5)
where $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ denotes the group of symmetry transformations preserving the Hamiltonian, and $\textnormal{U}_{\textnormal{TPS}}$ denotes the group of unitary transformations preserving the TPS.
To see if the necessary condition (5) can be satisfied, we can compare the dimensions of these two groups or of some of their subgroups.
The TPS is invariant under unitary transformations of $\mathcal{H}$ acting restrictively on each of the factor Hilbert spaces $\mathcal{H}_{j}$, and under the permutations of equal dimension factors $\mathcal{H}_{j}$ [11, 3]. The group of local unitary transformations acting only on $\mathcal{H}_{j}$ is $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}$. But since $\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}\cap\textnormal{U}\big{(}\mathcal{H}_{k}\big{)}=\textnormal{U}(1)$ for $j\neq k$, local phase changes are equivalent to the global ones. Taking this (and the fact that permutations of the factors don’t affect the dimension) into account, it follows that
$\dim\textnormal{U}_{\textnormal{TPS}}$ $\leq\sum_{j=1}^{n}\big{(}\dim\textnormal{U}\big{(}\mathcal{H}_{j}\big{)}-\underbrace{\dim\textnormal{U}(1)}_{\textnormal{local}}\big{)}+\underbrace{\dim\textnormal{U}(1)}_{\textnormal{global}}\big{)}$ (6)
$=\sum_{j=1}^{n}d_{j}^{2}-n+1,$
where $d_{j}:=\dim\mathcal{H}_{j}$, for all $j$.
Since the dimension of the Hilbert space $\mathcal{H}$ is $\dim\mathcal{H}=d_{1}\cdot\ldots\cdot d_{n}$, the dimension of the group of symmetry transformations preserving the Hamiltonian satisfies
$d_{1}\cdot\ldots\cdot d_{n}\leq\dim\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}\leq d_{1}^{2}\cdot\ldots\cdot d_{n}^{2},$ (7)
The lower bound is attained if all eigenvalues of $\widehat{\mathbf{H}}$ are distinct, and the upper bound is attained if all eigenvalues are equal.
If the necessary condition (5) were satisfied, any maximal commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ would be included in a maximal commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$. I will show that this doesn’t happen.
The dimension of any maximal commutative subgroup of $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ is at most
$\textnormal{D}_{\widehat{\mathbf{H}}}=d_{1}\cdot\ldots\cdot d_{n}.$ (8)
The dimension of any maximal commutative subgroup of $\textnormal{U}_{\textnormal{TPS}}$ is at most
$\textnormal{D}_{\textnormal{TPS}}=\sum_{j=1}^{n}(d_{j}-1)+1=\sum_{j=1}^{n}d_{j}-n+1.$ (9)
Since $d_{j}\geq 2$ for all $j$, for all $n>1$ we have
$d_{1}\cdot\ldots\cdot d_{n}>\sum_{j=1}^{n}d_{j}-n+1.$ (10)
Then, the necessary condition (5) can’t be satisfied, and there are symmetries preserving the Hamiltonian that don’t preserve the TPS. Therefore, the TPS can’t be uniquely determined by the Hamiltonian’s spectrum.
To obtain a unique TPS, additional parameters need to be fixed, and they can’t come from the spectrum.
Let’s see how fast does the number of additional parameters needed grows as $n$ grows. If all subsystems are qubits or qudits of the same dimension $d_{j}=d$, the dimensions of the two groups $\textnormal{U}_{\textnormal{TPS}}$ and $\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ are
$\begin{array}[]{lll}\dim&\textnormal{U}_{\textnormal{TPS}}&=1+n(d^{2}-1),\\ \dim&\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}&\geq d^{n}.\\ \end{array}$ (11)
We see that $\dim\textnormal{U}_{\textnormal{TPS}}$ is linear in $n$, while $\dim\textnormal{U}\big{(}\widehat{\mathbf{H}}\big{)}$ is exponential in $n$. Therefore, the number of additional continuous parameters needed to fix the TPS uniquely is exponential in the number of qudits. ∎
III Discussion: TPS, entanglement, and emergent spacetime
Now we’ll see that, even if the math would have allowed some Hamiltonians to determine a unique TPS, such a Hamiltonian would not generate a unitary dynamics able to entangle subsystems. Of course, we now know from Theorem 1 that the TPS can’t be unique, but it’s interesting to explore some physical reasons why this idea couldn’t work to begin with. This will also help seeing why we can’t declare the resulting TPSs to be equivalent in any physically meaningful way.
The operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$, $t\in\mathbb{R}$, commute with $\widehat{\mathbf{H}}$, so if the TPS is unique, when they are used as symmetry transformations, they should preserve it. But they also act as evolution operators of the system,
$\widehat{\mathbf{\rho}}(t)=e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}\widehat{\mathbf{\rho}}(0)e^{\frac{i}{\hbar}\widehat{\mathbf{H}}t}.$ (12)
Consider a state that is initially separable,
$\widehat{\mathbf{\rho}}(0)=\widehat{\mathbf{\rho}}_{1}(0)\otimes\widehat{\mathbf{\rho}}_{2}(0)\otimes\ldots.$ (13)
A state of the form $\ldots\otimes\widehat{\mathbf{1}}\otimes\widehat{\mathbf{\rho}}_{j}(0)\otimes\widehat{\mathbf{1}}\otimes\ldots$ is from the algebra of observables $\mathcal{A}_{j}$. From equation (4) it follows that it evolves into a state of the form
$\ldots\otimes\widehat{\mathbf{1}}\otimes\widehat{\mathbf{\rho}}^{\prime}_{\sigma(j)}(t)\otimes\widehat{\mathbf{1}}\otimes\ldots.$ (14)
Since $\widehat{\mathbf{\rho}}(t)$ is continuous in $t$, $\widehat{\mathbf{\rho}}^{\prime}_{\sigma(j)}(t)$ must retain its position $j$, so $\sigma(j)=j$ for all $j$. From equation (14) and the first condition in (2),
$\widehat{\mathbf{\rho}}(t)=\widehat{\mathbf{\rho}}^{\prime}_{1}(t)\otimes\widehat{\mathbf{\rho}}^{\prime}_{2}(t)\otimes\ldots.$ (15)
Therefore, such a unitary evolution can’t entangle the subsystems determined by the unique TPS. But in the real world, unitary evolution can entangle subsystems. Therefore, regardless of the method we use to get a unique TPS only from the spectrum, such a TPS can’t be the TPS of systems that can change their entanglement.
In addition, the entropy of entanglement is invariant under the symmetry transformations from $\textnormal{U}_{\textnormal{TPS}}$. Therefore, two TPSs determining different entanglement entropies should be nonequivalent. Since for unitary evolution the entropy of entanglement is continuous in time, symmetry transformations of the form $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ that change the entropy of entanglement can’t be from the group $\textnormal{U}_{\textnormal{TPS}}$. They transform the TPS into infinitely many nonequivalent TPSs resulting from the spectrum by the same method.
One may still hope that the TPS, while not unique, it is unique up to non-physical symmetry transformations. For example that non-uniqueness can be removed by taking equivalence classes of TPSs. But then again, the entanglement properties of such a system would not be able to change.
One may be distracted by using the operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ to prove that a unique TPS prevents entanglement, and propose that the TPS is determined up to the one-parameter group of symmetries $\big{(}e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}\big{)}_{t\in\mathbb{R}}$. This may even seem like a nice feature, because then the method would seem to recover not something about the system at a fixed $t$, but something about its entire history. But this would be a red herring: I only chose the operators $e^{-\frac{i}{\hbar}\widehat{\mathbf{H}}t}$ as an example, a useful one since we know that the entanglement can be changed by unitary time evolution. But in fact there are many other one-parameter groups that commute with $\widehat{\mathbf{H}}$, each of them being able to generate TPSs with different entanglement invariants, and therefore they can’t be equivalent.
It was proposed that, if a TPS emerges uniquely from the spectrum, the factors of the TPS can be identified with the points of the space, or whatever discrete building blocks of space we think should replace points in a quantum theory of gravity [2, 1].
The idea to recover the continuous $3$d-space from discrete building blocks is clever and it makes sense [5], and indeed these building blocks correspond to tensor product structures.
Then, if the TPS were unique, the distances could be recovered by using information that is not contained in the spectrum of the Hamiltonian, but that can be extracted from the state $\widehat{\mathbf{\rho}}$ and its relation with the TPS. Then maybe the $3$d space manifold structure would emerge at large scales. Carroll and Singh use the mutual information for this [2, 1].
But now we’ve seen that even if the resulting TPS were unique, unitary evolution wouldn’t be able to change the entanglement invariants of the TPS. The resulting spacetime metric, defined by using the mutual information, would be degenerate in the time direction, so the resulting geometry would be Carrollian (“Carrollian spaces” are named after Lewis Carroll [7, 4]). Everything would be frozen in time, and space would be absolute.
But more importantly, if the building blocks of space correspond to the factors of the TPS, Theorem 1 implies that they can’t emerge uniquely from the spectrum. Additional observables (along with their physical meaning [9]) are needed to begin with.